OpenStudy (anonymous):

A square based, box-shaped shipping crate is designed to have a volume of 16 ft^3. The material used to make the base costs twice as much (per ft^2) as the material in the sides, and the material used to make the top costs half as much (per ft^2) as the material in the sides. What are the dimension of the crate that minimize the cost of materials?

OpenStudy (anonymous):

@amistre64 please help I'm struggling

OpenStudy (anonymous):

@robtobey

OpenStudy (amistre64):

A square based, box-shaped shipping crate is designed to have a volume of 16 ft^3. b*b*h = 16 The material used to make the base costs twice as much (per ft^2) as the material in the sides, 4sides of area bh 1base of area bb and the material used to make the top costs half as much (per ft^2) as the material in the sides. 1top of area bb Total surface area?: 4bh + 2bb hmmm

OpenStudy (amistre64):

its a rate of change problem i can see that

OpenStudy (anonymous):

so far i have got to exactly where you are at

OpenStudy (amistre64):

We need a total Cost equation to minimise

OpenStudy (amistre64):

lets say S = cost for sides B = cost of base = 2S T = cost of top = S/2

OpenStudy (amistre64):

Cost = 4bhS + bb2S + bbS/2

OpenStudy (amistre64):

that should be doable, but a gotta go ..... for a min :)

OpenStudy (anonymous):

ok I'll keep trying to figure this out

OpenStudy (amistre64):

lets implicit this, and recall that the S is a constant price

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

Cost = 4S bh + 2S b^2 + S b^2/2 Cost' = 4S(b'h+bh') + 2S(2b b') + S(2b b')/2 Cost' = 4S(b'h+bh') + 4S bb' + S bb' Recall that we are given: V = bbh V' = 2bb' h+b^2 h' ; and since the volume remains constant, V' = 0 0 = 2bh b'+b^2 h' -2bh b' = b^2 h' -2b' h = bh'; might be able to use this to simplify stuff To minimize costs, we want Cost' to be 0 0 = 4S(b'h+bh') + 4S bb' + S bb' 0 = 4Sb'h + 4Sbh' + 5S bb' -4S b'h = 4Sbh' + 5S bb' ; -4b' h = 2bh' 2S bh' = 4Sbh' + 5S bb' ; divide out the S 2bh' = 4bh' + 5 bb' -2bh' = 5 bb' ; divide of the b -2h' = 5b' looks to be getting someplace .... what was the question again?

OpenStudy (amistre64):

oh yeah, find bh that minimize costs

OpenStudy (amistre64):

h' = -5b'/2 h = -5b/2 .... gotta think if thats useful or not now :)

OpenStudy (amistre64):

if we let b' = 1 ..... 0 = 4S(b'h+bh') + 4S bb' + S bb' 0 = 4S(h+bh') + 4S b + S b and h' = -5/2, while h=-5b/2 0 = 4(-5b/2-5b/2) + 4b + b 0 = 4(-5b) + 4b + b 0 = b(-20+4+1) grrr, missing something im sure

OpenStudy (amistre64):

0 = 4(-5b/2-5b^2/2) + 4b + b ^^ 5b^2 , not 5b ;)

OpenStudy (amistre64):

0 = -20b/2-20b^2/2 + 5b 0 = -10b - 10b^2 + 5b 0 = -10b^2 -5b 0 = 5b(2b + 1) b = 0, or b = -1/2; lol .... i wonder if 1/2 is suitable

OpenStudy (anonymous):

says that the dimensions should be \[\frac{ 4 }{ \sqrt[3]{5} },\frac{ 4 }{ \sqrt[3]{5} },5^{\frac{ 2 }{ 3 }}\]

OpenStudy (amistre64):

thats the answer key?

OpenStudy (anonymous):

yessir

OpenStudy (amistre64):

on paper i was getting something like b = 4/sqrt(2) :)

OpenStudy (amistre64):

we can assume the cost to be an arbitrary value like 1 for simplicity

OpenStudy (anonymous):

well that's extremely close. Much closer than anything i would've gotten without your help

OpenStudy (amistre64):

Cost = 1(4bh) + 2(b^2) + b^2/2 sides base top Cost = 4bh + 5b^2/2 is a simpler cost function to me 16 = b^2h, such that h=16/b^2

OpenStudy (anonymous):

that's what i started out doing

OpenStudy (amistre64):

Cost = 4b16/b^2 + 5b^2/2 Cost = 64b^(-1) + 5/2 b^2 Cost' = -64b^(-2) + 5b ; min max is when Cost' = 0 5b = 64b^-2 5b^3 = 64 b = cbrt(64/5)

OpenStudy (amistre64):

not real sure why i didnt simplify it beforehand :)

OpenStudy (amistre64):

\[b=\sqrt[3]{\frac{64}{5}}=\frac{4}{\sqrt[3]{5}}\]

OpenStudy (anonymous):

how did you get the -64b^-2

OpenStudy (amistre64):

do you agree that the cost function is: Cost = 4bh + 5b^2/2

OpenStudy (amistre64):

do you agree that the constraint for the volume is: 16 = b^2h, giving us h = 16b^(-2) ?

OpenStudy (anonymous):

oh yea

OpenStudy (amistre64):

if so then: Cost = 4bh + 5b^2/2 Cost = 4b(16b^(-2)) + 5b^2/2 Cost = 64b^(-1) + 5b^2/2

OpenStudy (amistre64):

the rest is just a power derivative

OpenStudy (anonymous):

ok see i didn't realize we were taking the derivative yet

OpenStudy (amistre64):

had to take it at sometime ;) but yeah, it worked up simpler than i started .... getting old bites ;)

OpenStudy (anonymous):

lol. Hey is there a way to save conversations or something so i could come back to this later?

OpenStudy (anonymous):

A solution using Mathematica for the calculations is attached.

OpenStudy (anonymous):

thanks a lot man means alot that you took the time out to help as well